OpenStudy (anonymous):

The quotient of a number and five is eight less than the number. What is the number?

OpenStudy (kropot72):

Let the number be x. \[\frac{x}{5}+8=x\] \[x-\frac{x}{5}=8\] \[\frac{4x}{5}=8\] Can you solve it now?

OpenStudy (kropot72):

Multiply both sides by 5: 4x = 40 x = 10